Calculate the position of the planet 10^5 kg which is placed between the planets of 10^20kg and 10^24 kg which are placed at 6000km apart hence it will get zero gravitational force.(I have solved the question,I only want the working and Correct and complete answer) 

Grade:9

1 Answers

Vikas TU
4 years ago
Hiii 
Gravitational force = GM1*M2/d2
Let us assume distance of 10^5 Kg particles from 10^20 Kg be x 
So gravitation force between them = G *10^25 /x2
Gravitational force between 10^5 and 10^24 kg mass will be = G* 10^29 /(6000-x)2
It is getting a zero gravitational force so, G*10^25 /x2 = G* 10^29 /(6000-x)2
(6000 – x)2  = 10^4 * x2
solve for x and apply quadratic equation formula , 
You will get X = 59.40 Km